OpenStudy (anonymous):

Describe an infinite geometric series with a beginning value of 2 that converges to 10. What are the first 4 terms of the sequence? Please Help

OpenStudy (anonymous):

Ok cool Thanks :)

alones (alones):

I hope it would help <3

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

ok sorry i had to double check the math before i answered so you have a geometric series: \[2,2r,2r^2 ... 2r^n\] where \[\lim_{n \rightarrow \infty} r^n = 5\] Note another limit used for continuous compounding, how "e" is defined \[\rightarrow \lim_{n \rightarrow \infty} (1+\frac{i}{n})^n = e^i\] Now if you let : \[r = 1+\frac{i}{n} , i = \ln(5)\] Then \[\lim_{n \rightarrow \infty}r^n = (1+\frac{\ln(5)}{n})^n = e^{\ln 5} = 5\]

OpenStudy (anonymous):

Whoa Thanks :)

OpenStudy (dumbcow):

yw